OpenStudy (anonymous):

what is lcm for 1,2,3,4,5,6,7,8,9,10 and show how you get the answer

OpenStudy (shadowfiend):

We can try to do some prime factorization. The highest prime in this list is 7, so we look at each number in terms of its prime factors: \begin{align} 1 &= 2^0 \times 3^0 \times 5^0 \times 7^0\\ 2 &= 2^1 \times 3^0 \times 5^0 \times 7^0\\ 3 &= 2^0 \times 3^1 \times 5^0 \times 7^0\\ 4 &= 2^2 \times 3^0 \times 5^0 \times 7^0\\ 5 &= 2^0 \times 3^0 \times 5^1 \times 7^0\\ 6 &= 2^1 \times 3^1 \times 5^0 \times 7^0\\ 7 &= 2^0 \times 3^0 \times 5^0 \times 7^0\\ 8 &= 2^3 \times 3^0 \times 5^0 \times 7^0\\ 9 &= 2^0 \times 3^2 \times 5^0 \times 7^0\\ 10 &= 2^1 \times 3^0 \times 5^1 \times 7^0 \end{align}

hero (hero):

Yeah, I started with 7 too. 840 is divisible by everything except 9

OpenStudy (shadowfiend):

To find the LCM, you can then multiply the highest power of each prime. In this case, we have \(2^3\), \(3^2\), \(5^1\), and \(7^1\): \[2^3 \times 3^2 \times 5^1 \times 7^1 = 6 \times 9 \times 5 \times 7 = 1890\] If I'm not mistaken.

hero (hero):

Nice

OpenStudy (shadowfiend):

Hah. Whoops. In my original list, I forgot to list \(7^1\) as the correct value for 7. Fortunately, I didn't make that mistake when I multiplied things out :)

OpenStudy (anonymous):

the answer is 2520 but i dont know how to get it

hero (hero):

Yeah, 1890 isn't divisible by 8

hero (hero):

I like your method shadow. I hope you can find your mistake

OpenStudy (shadowfiend):

Lol. Blah O_O

OpenStudy (anonymous):

2^ 3 ×3^ 2 ×5 ×7 =8×9×5×7=2520

OpenStudy (anonymous):

2^3=8

OpenStudy (shadowfiend):

Wait, you mean 2^3 isn't 6? Lol. Thanks. Epic failure on my part.

hero (hero):

:D

hero (hero):

So it does work. I love it

hero (hero):

Shadow deserves the medal for this one. He made a simple mistake, but provided the technique. Marina also deserves one for correcting the mistake

OpenStudy (anonymous):

I agree. shadow did most of the work and explanation.